OpenStudy (anonymous):

x+y=7 can someone help me I will medal

OpenStudy (lifeisadangerousgame):

What are you trying to do? Solve for x or y?

OpenStudy (anonymous):

x

OpenStudy (lifeisadangerousgame):

Alright, you need to get x by itself on the left side. What is the first step do you think?

OpenStudy (anonymous):

add x to both sides

OpenStudy (lifeisadangerousgame):

No. You want x to stay on the left side, but y is there too. How do we get y on the right side?

OpenStudy (anonymous):

add 7 to both

OpenStudy (lifeisadangerousgame):

No, that will bring the seven onto the left side. You want the seven AND the y to be on the right side, what could you do to make 'y' go on the right side of the equals sign?

OpenStudy (anonymous):

idk how

OpenStudy (lifeisadangerousgame):

Subtract y from both sides. This is algebra. x + y = 7 -y -y x = y -7 Does that make sense?

OpenStudy (anonymous):

yes thx